OpenStudy (anonymous):

what is \[\LARGE \int e^{1 + \ln x}dx\] i just need the step-by-step solution...no need for answers ^_^

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

That is an awesome integral!!!

myininaya (myininaya):

\[\text{ clue" } e^{x+y}=e^xe^y\]

myininaya (myininaya):

Law of exponents to the rescue :)

myininaya (myininaya):

@mr.awesome any questions?

OpenStudy (anonymous):

where is he?

myininaya (myininaya):

he ran away :(

OpenStudy (anonymous):

phooey....:(

OpenStudy (anonymous):

sorry my momma called (telephone) can you guys guide me more? is it \[\huge \int e^1e^{\ln x}?\]

myininaya (myininaya):

e is a constant pull that sucka out and guess what e^(ln(x)) = ____ for x>0

OpenStudy (anonymous):

naw... leave it in....

myininaya (myininaya):

Remember exponential function and the natural log function are inverse functions

myininaya (myininaya):

\[e \int\limits_{}^{}e^{\ln(x)} dx\] you can leave the constant multiple in there if you want

OpenStudy (anonymous):

@dpaInc what happens if you leave e in there???

myininaya (myininaya):

but what does \[e^{\ln(x)}=?, x>0\]

OpenStudy (anonymous):

i dunno can u write it in there?

OpenStudy (anonymous):

\[\huge \int ee^{\ln x}??\]

myininaya (myininaya):

ok if you answer my question we can get even further

OpenStudy (anonymous):

uhh sorry..i was amused by dpainc's suggestion...what were you saying?

OpenStudy (dumbcow):

i think where they were headed is \[e^{\ln x} = x\] \[\rightarrow e \int\limits_{}^{} x dx = \frac{e}{2} x^{2} +C\]

myininaya (myininaya):

right I was trying to get mr.awesome to fill in that blank